# skipabove=5pt, innertopmargin=0pt

### v1

@@ 5 Experiments
We compare our algorithm with state-of-the-art methods for jailbreaking in both the token-space and the prompt-space. Specifically, we include GCG [ZWC+23], PAIR [CRD+24], TAP-T [MZK+24], and AutoDAN-turbo [LLS+24b] which is an extension of AutoDAN [LXCX24]. Additionally, we incorporate results from [ACF24], even though some of their methods go beyond just modifying the target LLM’s input and employ pre-filling attacks or alter the target model’s system prompt. As we limit our comparison to methods that interact with the target LLM through its input, we use only their crafted template along with random search, which we refer to as “Prompt + Random Search”. We use Attack Success Rate (ASR) as the main metric for comparison. We execute our algorithm against some of the safest LLMs according to [Swa] leaderboard, including both open-source (white-box) and proprietary (black-box) models. The HarmBench judge [MPY+24] is deployed to evaluate the target LLM’s responses due to its high alignment with human evaluations [SLB+24]. We test our algorithm on 50 uniformly sampled tasks selected from standard behaviors in the Harmbench dataset [MPY+24]. We manually verify all proposed jailbreaks to avoid false positives.
@@
Attacker models As for the attacker model, we use LLMs without any moderation mechanism to ensure compliance. Specifically, we use “Vicuna-13b-v1.5” (Vicuna) [CLL+23] and “Mixtral-8x7B-v0.1” (Mixtral) [AI24a]. The details of our curated system prompt for the attacker are given in Appendix C. The temperature of the attacker LLM is set to 1.0, as exploration is critical in our framework.
Feedback LLM and Refiner LLM. For a fair comparison with other attacker-based methods, we use the same model for the Feedback LLM, Refiner LLM, and attacker model, as they are all part of the attacking team. This setup isolates the effect of each method from differences in model capability. Details of their configuration and the rationale behind these choices are in Appendix C. At each call of Feedback LLM, we divide the $n$ sorted attacking prompts into $k$ buckets and uniformly sample one prompt from each bucket. Feedback LLM then evaluates only $n/k$ prompts at a time. This strategy facilitates exploration in the search algorithm by increasing the diversity of feedback. While the prompt with the lowest loss is more likely to succeed in jailbreaking, comparing the other prompts can provide more informative feedback and enhance the overall effectiveness of the optimization process.
@@ 5.1 Attack Success Rate
For all methods that rely on an attacker LLM to generate the attacking prompts—except for AutoDAN-Turbo [^1] We were unable to reproduce results with AutoDAN-Turbo’s official implementation; the attacker model did not generate the expected outputs even after substituting Llama-3-70 with Mixtral. we have deployed Mixtral [^2] Note that the standard versions of PAIR and TAP-T use Vicuna. The use of a stronger model such as Mixtral here leads to a higher ASR than their reported results.. The main results are presented in Table 1. Except for Llama-3-8B, our method achieves the best ASR among both the token-space and the prompt-space attempts. This is significant because token-space algorithms operate without constraints to be semantically meaningful, making it easier to find a jailbreaking prompt. However, as shown in Table 1, for target models that have been adversarially trained against token-level jailbreaking such as Llama-3-8B-RR [ZPW+24] and R2D2 [MPY+24], these algorithms largely fail since as the rely on eliciting only a handful of tokens. In Appendix D, we provide jailbreak examples and details of how our algorithm works w.r.t. those examples in Figures 10 and 11.
@@
Target model GCG Prompt + Random Search AutoDAN-Turbo PAIR TAP-T Adversarial Reasoning
@@
Llama-3-8B 44% 100% 62% 66% 76% 88%
Llama-3-8B-RR 2% 0% - 22% 32% 44%
Mistral-7B-v2-RR 6% 0% - 32% 40% 70%
R2D2 0% 12% 84% 98% 100% 100%
@@
The ASR of all the algorithms that rely on an LLM for generating the attacking prompts varies by the capabilities of that LLM. These capabilities, however, can be improved by scaling the test-time computation [SLXK24]. We demonstrate the efficacy of our algorithm by using a weaker attacker model such as Vicuna. We compare the ASR of our algorithm to those of PAIR [CRD+24] and TAP-T [MZK+24], all targeting Llama-3-8B. As shown in Table 2, our algorithm achieves an ASR of $64\%$ with Vicuna—more than three times the ASR achieved by PAIR and TAP-T. Notably, it nearly attains the same ASR as PAIR with Mixtral as the attacker– a much stronger LLM than Vicuna. Furthermore, our method uses a very simple system prompt for the attacker LLM (see Appendix C) compared to methods such as PAIR and TAP-T. This highlights the effectiveness of optimally scaling test-time computation rather than scaling the model. This aligns with a broader trend in the reasoning literature [SLXK24].
@@ 5.2 Multi-shot transfer attacks
Given the infeasibility of obtaining the log-prob vectors in black-box models, we evaluate the success of our algorithm using two transfer methods. We perform the transfer by optimizing the loss function on a surrogate white-box model and then applying the derived adversarial prompt to the target black-box model. A common approach is to transfer the prompt that jailbreaks or yields the lowest loss on the surrogate model [ZWC+23]—we call this a “one-shot” transfer attack. However, this does not always result in an effective attack as the loss function serves only as a heuristic in the transfer. We improve effectiveness by using a scheme that queries the target model with all the attacking prompts collected from executing the algorithm ($n$ prompts per iteration). We call this a “multi-shot” transfer. We show that the transfer success significantly increases with this scheme. We use the loss values from three white-box models: Llama-2, Llama-3-RR, and R2D2 as the surrogate loss in our algorithm in Section 4 to conduct attacks on black-box models. [ZWC+23] demonstrates that aggregating losses from multiple target models enhances the transferability of the final prompt compared to relying on a single model. If we have surrogate models $\mathbb{M}_{1},\dots,\mathbb{M}_{r}$, the aggregated loss function is $\frac{1}{r}\sum_{i=1}^{r}\mathcal{L}_{\mathbb{M}_{i}}(\mathbf{y}_{I},P)$, where each loss is calculated according to Section 3. We run Algorithm 1 with this loss to evaluate the attacking prompts, and to assign the scores in Equation 4.3. We assess the effectiveness of using the aggregated loss as the surrogate for $r=3$ using the mentioned above models. Details of our transfer method using $r$ models for surrogate loss estimation are presented in Algorithm 2.
@@
2:Run Algorithm 1 with $\text{Run Algorithm~{}1 with }\frac{1}{r}\sum_{i=1}^{r}\mathcal{L}_{\mathbb{M}_{i}}(\mathbf{y}_{I},P)$
@@
Table 3 presents the results of our experiments in comparing “one-shot” and “multi-shot” settings. The multi-shot approach significantly improves the ASR across all tested models, with the exception of Cygnet [ZPW+24]. Cygnet, a variant of Llama-3-RR, incorporates a strict safety filter that blocked almost all of our attempts. Indeed, the success rate of any other jailbreaking algorithm is near zero for Cygnet [Swa]; therefore, our findings demonstrate a higher ASR compared to existing literature. Notably, we achieve 94% on Llama-3.1-405B (without Llama Guard) [AI@24b], 94% on GPT4o [Ope24b], and 66% on Gemini-1.5-pro. For models with stricter moderation mechanisms, such as OpenAI o1-preview and Claude-3.5-Sonnet, using the average loss boosts the ASR. For instance, [Ope24b] reports an ASR of 16% for the o1-preview model. Using our attack, this increases to 56%. Further details about the transfer experimental setup are provided in Appendix A. Also, fig. 6 presents the vulnerabilities of different models across the six categories in Harmbench.
Is DeepSeek a safe model? We evaluated our method on the recently released DeepSeek model [DA25] using the aggregated surrogate loss. Our approach achieved 100% ASR, indicating that DeepSeek failed to block a single adversarial reasoning attack. This outcome suggests that DeepSeek’s cost-efficient training methods may have compromised its safety mechanisms. Compared to other frontier models, it lacks robust guardrails, making it highly susceptible to algorithmic jailbreaking. Due to time and resource constraints, we did not evaluate other baselines on this model and, therefore, excluded DeepSeek from Table 3.
@@ 5.3 Ablation studies
We investigate whether our method fulfills its objectives in (i) minimizing the loss function; and (ii) generating the feedback based on the ordered attacking prompts. Besides, we demonstrate that our method finds more jailbreaks in later iterations compared to previous work.
@@
Our algorithm aims to optimize a loss function defined over a string. To assess its effectiveness, we can probe $\mathbb{V}(S)$ in Equation 4.3 over iterations. Figure 3 illustrates this loss progression for Llama-2-7b, Llama-3-8b, Llama-3-8b-RR, and R2D2, with Mixtral as the attacker and averaged over 50 tasks. The results showcases a quantitative decrease in the minimum loss for all target models until approximately the 10th iteration, after which the loss exhibits slight oscillations. Notably, for R2D2, the loss converges to zero despite the safety measures. As depicted in Figure 3, the loss curve for Llama-3-8b-RR shows a significant gap compared to what its fine-tuned from, Llama-3-8b, and despite this gap, our algorithm achieves $44\%$ of ASR. To investigate this, we have plotted the loss separately for successful and failed jailbreaking attempts in Figure 3. This figure shows that, on average, successful attempts start and end with lower loss values than failures. This demonstrates the utility of Section 3 as a heuristic; Although the absolute loss value may remain high, its relative value serves as an informative metric. Consequently, comparing the losses of multiple attacking prompts provides feedback that guides the algorithm toward more effective attempts. Unsurprisingly, the jailbreaks do not begin with their desired output string $\mathbf{y}_{I}$ for Llama-3-8b-RR. However, in Appendix D, we show that there are other possibilities such as initially refusing followed by compliance. This helps us establish our work as an attack in the prompt space that rigorously optimizes a loss function, similar to algorithms in the token-space, but with significantly fewer number of iterations. In contrast, the number of iterations for the token-space algorithms can be as high as $10^{4}$ [ACF24].
Figure 3: left: We plot the objective of our optimization (Equation 4.3) over iteration to demonstrate that the refinement process is effective. right: Shows that despite the high value of the loss, its relative value still provides a signal for identifying the successful attempts, and hance, a heuristic for our method.
@@
$\frac{\mathbb{P}_{\mathbb{A}}\big{(}P_{a}|\mathbb{R}(S,F)\big{)}}{\mathbb{P}_{\mathbb{A}}{\big{(}P_{b}|\mathbb{R}(S,F)\big{)}}}\geq\frac{\mathbb{P}_{\mathbb{A}}(P_{a}|S)}{\mathbb{P}_{\mathbb{A}}{(P_{b}|S)}}\text{ if}\ \ \mathcal{L}_{\mathbb{T}}(\mathbf{y}_{I},P_{a})\leq\mathcal{L}_{\mathbb{T}}(\mathbf{y}_{I},P_{b})\text{ and}\ a,b\in\{1,\cdots,n\}$ (5.1)
Here, $\mathbb{P}_{\mathbb{A}}(P|S)$ denotes the probability of the attacker LLM generating prompt $P$ conditioned on the reasoning string $S$. I.e., $\mathbb{P}_{\mathbb{A}}(P|S)=\prod_{i=1}^{l}\mathbb{P}_{\mathbb{A}}\big{(}p_{i}|[S,p_{1:i-1}]\big{)}$ where $P=[p_{1},\cdots,p_{l}]$. To evaluate this condition, we analyze the generation probabilities for a set of prompts before and after applying feedback to the reasoning string. When the attacker is Vicuna and the target is Llama-3-8B-RR, for 10 random tasks each with 10 iterations (totaling 100 function calls), we compute the difference in Cross-Entropy of the prompts conditioned on the original and updated reasoning strings: $-\log\Big{(}\mathbb{P}_{\mathbb{A}}\big{(}P_{a}|\mathbb{R}(S,F)\big{)}\Big{)}+\log\Big{(}\mathbb{P}_{\mathbb{A}}(P_{a}|S)\Big{)}$. If $a$ denotes the ordered index, this function must be increasing with respect to $a$ in order to satisfy Equation 5.1. Figure 4 presents the average results across all calls for $n=16$, showing that at each iteration, the curve roughly increases, with a slight decline in the last two prompts. This decline can be attributed to the "Lost in the Middle" effect, where models tend to focus more on the most recent text [LLH+23]. Nonetheless, the value of the last two prompts do not fall below the value of the initial prompts, indicating that the generated feedback remains meaningful and not misleading. In Appendix B, We conduct two more experiments to shed more light on the functionality of Feedback LLM. As a sanity check, we demonstrate that the reversing the order of the given attacking prompts to the Feedback LLM leads to semantically reversed feedbacks, and a drop in the ASR due to moving in the wrong direction in the prompt-space.
@@
The standard version of PAIR runs for only 3 iterations. We increased this number to $T=15$, matching it with our algorithm. Both algorithms use Mixtral to attack Llama-2-7B. Figures 5(a) and 5(b) depict the number of tasks are jailbroken at each iteration for PAIR and our algorithm, respectively. PAIR achieved only two additional successful jailbreaks after the third iteration, accounting for 11% of successful attempts. In contrast, our algorithm accomplished 37% of jailbreaks after the third iteration, demonstrating a more effective utilization of iterations. Distributions of jailbreaks for other models are presented in Figures 8 and 9. As anticipated, a higher percentage of jailbreaks occur in later iterations for models with stricter safety measures that necessitate more extensive search; for instance, 57% after the third iteration for Llama-3-8B-RR (Figure 8(a)). PAIR relies on the reasoning capabilities of the attacker LLM to modify subsequent attacking prompts. Therefore, after encountering a few refusals from the target model, PAIR tends to deviate from the original intent to avoid the refusals. A couple of examples of this phenomenon are presented in Appendix D.
6 Acknowledgment
This work of HH and GP was supported by the NSF Institute for CORE Emerging Methods in Data Science (EnCORE). SM was supported by both EnCORE and RobustIntelligence. The authors wish to thank Alexander Robey for helpful discussions.

### v2

@@ 5 Experiments
We compare our algorithm with state-of-the-art methods for jailbreaking in both the token-space and the semantic-space. Specifically, we include GCG (Zou et al., 2023), PAIR (Chao et al., 2024b), TAP-T (Mehrotra et al., 2024), Rainbow Teaming (Samvelyan et al., 2024), and AutoDAN-turbo (Liu et al., 2024a) which is an extension of AutoDAN (Liu et al., 2024b). Additionally, we incorporate results from (Andriushchenko et al., 2024a), even though some of their methods go beyond just modifying the target LLM’s input and employ pre-filling attacks or alter the target model’s system prompt. As we limit our comparison to methods that interact with the target LLM through its input, we use only their crafted template along with random search, which we refer to as “Adaptive Attack”. We use Attack Success Rate (ASR) as the main metric for comparison. We execute our algorithm against some of the safest LLMs including both open-source (white-box) and proprietary (black-box) models. The HarmBench judge (Mazeika et al., 2024) is deployed to evaluate the target LLM’s responses due to its high alignment with human evaluations (Souly et al., 2024; Chao et al., 2024a). We test our algorithm on 50 uniformly sampled tasks selected from standard behaviors in the Harmbench dataset (Mazeika et al., 2024). We manually verify all proposed jailbreaks to avoid false positives. (Read Appendix B for more details about verification.)
@@
Attacker models As for the attacker model, we use LLMs without any moderation mechanism to ensure compliance. Specifically, we use “Vicuna-13b-v1.5” (Vicuna) and “Mixtral-8x7B-v0.1” (Mixtral) (Mistral AI, 2024). The details of our curated system prompt for the attacker are given in Appendix D. The temperature of the attacker LLM is set to 1.0, as exploration is critical in our framework.
Feedback LLM and Refiner LLM. For a fair comparison with other attacker-based methods, we use the same model for the Feedback LLM, Refiner LLM, and attacker model, as they are all part of the attacking team. This setup isolates the effect of each method from differences in model capability. Details of their configuration and the rationale behind these choices are in Appendix D. At each call of Feedback LLM, we divide the $n$ sorted attacking prompts into $k$ buckets and uniformly sample one prompt from each bucket. Feedback LLM then evaluates only $n/k$ prompts at a time. This strategy facilitates exploration in the search algorithm by increasing the diversity of feedback. While the prompt with the lowest loss is more likely to succeed in jailbreaking, comparing the other prompts can provide more informative feedback and enhance the overall effectiveness of the optimization process.
@@ 5.1 Attack Success Rate
For all methods that rely on an attacker LLM to generate the attacking prompts we deploy Mixtral [^1] Note that the standard versions of PAIR and TAP-T use Vicuna. The use of a stronger model such as Mixtral here leads to a higher ASR than their reported results.. The main results are presented in Table 1. Except for Llama-3-8B, our method achieves the best ASR among both the token-space and the semantic-space attempts. Notably, token-space algorithms operate without constraints to be semantically meaningful, thereby facilitating the search for a jailbreaking prompt. However, as shown in Table 1, for target models that have been adversarially trained against token-level jailbreaking such as Llama-3-8B-RR (Zou et al., 2024) and R2D2 (Mazeika et al., 2024), these algorithms largely fail as they rely on eliciting only a few tokens. In Appendix E, we provide jailbreak examples and details of how our algorithm works w.r.t. those examples in Figures 10 and 11.
Additionally, in Table 2, we compare the average of overall queries per success among all the baselines. Although AutoDAN-Turbo and Rainbow Teaming require few evaluation-time prompts, they involve extensive upfront computation to identify initial strategies, so this should be incorporated into any comparison. While our method does require more queries than PAIR, it achieves higher success rates on difficult tasks that demand multiple refinement steps. In this manner, another useful comparison is the one provided in Figure 5 where we compare the number of jailbreaks at each iteration for both algorithms.
@@
Target model GCG Adaptive Attack Rainbow Teaming AutoDAN- Turbo PAIR TAP-T Adversarial Reasoning
@@
Llama-3-8B 44% 100% 26% 62% 66% 76% 88%
Llama-3-8B-RR 2% 0% 14% 26% 22% 32% 44%
Mistral-7B-v2-RR 6% 0% 30% 40% 32% 40% 70%
R2D2 0% 12% 70% 84% 98% 100% 100%
@@
GCG Adaptive Attack Rainbow Teaming AutoDAN- Turbo PAIR TAP-T Adversarial Reasoning
250 2600 6K 60K 33 20 48
Table 2: Comparison of number of queries-per-success for various methods. Average taken over five target models in Table 1.
@@
The ASR of all the algorithms that rely on an LLM for generating the attacking prompts varies by the capabilities of that LLM. These capabilities, however, can be improved by scaling the test-time computation (Snell et al., 2024). We demonstrate the efficacy of our algorithm by using a weaker attacker model such as Vicuna. We compare the ASR of our algorithm to those of PAIR (Chao et al., 2024b) and TAP-T (Mehrotra et al., 2024), all targeting Llama-3-8B. As shown in Table 3, our algorithm achieves an ASR of $64\%$ with Vicuna—more than three times the ASR achieved by PAIR and TAP-T. Notably, it nearly attains the same ASR as PAIR with Mixtral as the attacker– a much stronger LLM than Vicuna. Furthermore, our method uses a very simple system prompt for the attacker LLM (see Appendix D) compared to methods such as PAIR and TAP-T. This highlights the effectiveness of optimally scaling test-time computation rather than scaling the model. This aligns with a broader trend in the reasoning literature (Snell et al., 2024).
@@ 5.2 Multi-shot transfer attacks
Given the infeasibility of obtaining the log-prob vectors in black-box models, we evaluate the success of our algorithm using two transfer methods. We perform the transfer by optimizing the loss function on a surrogate white-box model and then applying the derived adversarial prompt to the target black-box model. A common approach is to transfer the prompt that jailbreaks or yields the lowest loss on the surrogate model (Zou et al., 2023)—we call this a “one-shot” transfer attack. However, this does not always result in an effective attack as the loss function serves only as a heuristic in the transfer. We improve effectiveness by using a scheme that queries the target model with all the attacking prompts collected from executing the algorithm ($n$ prompts per iteration). We call this a “multi-shot” transfer. We show that the transfer success significantly increases with this scheme. We use the loss values from three white-box models: Llama-2, Llama-3-RR, and R2D2 as the surrogate loss in our algorithm in Section 4 to conduct attacks on black-box models. (Zou et al., 2023) demonstrates that aggregating losses from multiple target models enhances the transferability of the final prompt compared to relying on a single model. If we have surrogate models $\mathbb{M}_{1},\dots,\mathbb{M}_{r}$, the aggregated loss function is $\frac{1}{r}\sum_{i=1}^{r}\mathcal{L}_{\mathbb{M}_{i}}(P,\mathbf{y}_{I})$, where each loss is calculated according to Section 3. We run Algorithm 1 with this loss to evaluate the attacking prompts, and to assign the scores in Equation 4.3. We assess the effectiveness of using the aggregated loss as the surrogate for $r=3$ using the mentioned above models. Details of our transfer method using $r$ models for surrogate loss estimation are presented in Algorithm 2.
@@
2:Run Algorithm 1 with $\text{Run Algorithm~{}1 with }\frac{1}{r}\sum_{i=1}^{r}\mathcal{L}_{\mathbb{M}_{i}}(P,\mathbf{y}_{I})$
@@
Table 4 presents the results of our experiments in comparing “one-shot” and “multi-shot” settings. The multi-shot approach significantly improves the ASR across all tested models, with the exception of Cygnet (Zou et al., 2024). Cygnet, a variant of Llama-3-RR, incorporates a strict safety filter that blocked almost all of our attempts. Indeed, the success rate of any other jailbreaking algorithm is near zero for Cygnet (Zou et al., 2024); therefore, our findings demonstrate a higher ASR compared to existing literature. Notably, we achieve 94% on Llama-3.1-405B (without Llama Guard) (Meta AI, 2024), 94% on GPT4o, and 66% on Gemini-1.5-pro. For models with stricter moderation mechanisms, such as OpenAI o1-preview and Claude-3.5-Sonnet, using the average loss boosts the ASR. For instance, (OpenAI, 2024b) reports an ASR of 16% for the o1-preview model. Using our attack, this increases to 56%. Further details about the transfer experimental setup are provided in Appendix B. Figure 6 presents the vulnerabilities of different models across the six categories in Harmbench.
DeepSeek-R1 evaluation We evaluate DeepSeek-R1 (DeepSeek-AI, 2025) using the aggregated surrogate loss. Our approach achieves 100% ASR, indicating that Deepseek fails to block a single adversarial reasoning attack. This outcome suggests that lacks robust guardrails, making it highly susceptible to algorithmic jailbreaking.
@@ 5.3 Ablation studies
We examine whether our method (i) minimizes the loss function; and (ii) relies on the feedback string for its effectiveness. Furthermore, we show that it uncovers more jailbreaks in later iterations than prior work.
@@
Our algorithm aims to optimize a loss function defined over a string. To assess its effectiveness, we can probe $\mathbb{V}(S)$ in Equation 4.3 over iterations. Figure 3 illustrates this loss progression for various targets with Mixtral as the attacker and averaged over 50 tasks. The results showcase a quantitative decrease in the minimum loss for all target models until approximately the 10th iteration, after which the loss exhibits slight oscillations. Notably, for R2D2, the loss converges to zero despite the safety measures. As depicted in Figure 3, the loss curve for Llama-3-8B-RR shows a significant gap compared to what its fine-tuned from, Llama-3-8B, and despite this gap, our algorithm achieves $44\%$ of ASR. To investigate this, we plot the loss separately for successful and failed jailbreaking attempts in Figure 3. This figure shows that, on average, successful attempts depict lower loss values than failures. This demonstrates the utility of Section 3 as a heuristic; Although the absolute loss value may remain high, its relative value serves as an informative metric. Unsurprisingly, the jailbreaks do not begin with their desired output string $\mathbf{y}_{I}$ for Llama-3-8B-RR. However, in Appendix E, we show that there are other possibilities such as initially refusing followed by compliance. Figure 3 helps us establish our work as an attack in the prompt space that effectively optimizes a loss, similar to algorithms in the token-space, but with significantly fewer number of iterations. The number of iterations for the token-space algorithms can be as high as $10^{4}$ (Andriushchenko et al., 2024a).
Figure 3: Up: We plot the objective of our optimization (Equation 4.3) over iteration to demonstrate that the refinement process is effective. Down: Shows that despite the high value of the loss, its relative value still provides a signal for identifying the successful attempts, and hance, a heuristic for our method.
Loss-guided reasoning is essential
. We explicitly construct the adversarial reasoning steps (see Figure 2) to ensure the relevance of such directions to decreasing the loss. To highlight the importance of this design and its distinction from relying solely on intrinsic CoT capabilities of LLMs, we conducted an experiment using PAIR with Deepseek-R1 as the attacker. As Table 5 shows, even advanced reasoning models such as Deepseek struggle when operating heuristically, and without our structured reasoning algorithm and external supervision via the loss. This aligns with recent work by Kritz et al. (2025, Fig. 6), where OpenAI-o3 underperforms as the attacker. Our conclusion is that finding the relevant directions by acquiring the loss-based feedback string is necessary to jailbreak stronger models.
Target model PAIR + Deepseek-R1 Adversarial Reasoning
Claude-3.5-Sonnet $16$% $36$%
o1-preview $16$% $56$%
Table 5: Comparison with PAIR when it is utilized with Deepseek, whereas our method uses Mixtral. This demonstrates that even use of strongest LLMs has limited impact on the ASR as long as they are not properly supervised and act heuristically instead.
@@
$\frac{\mathbb{P}_{\mathbb{A}}\big{(}P_{a}|\mathbb{R}(S,F)\big{)}}{\mathbb{P}_{\mathbb{A}}{\big{(}P_{b}|\mathbb{R}(S,F)\big{)}}}\geq\frac{\mathbb{P}_{\mathbb{A}}(P_{a}|S)}{\mathbb{P}_{\mathbb{A}}{(P_{b}|S)}}$ (5.1)
$\text{if}\ \ \mathcal{L}_{\mathbb{T}}(\mathbf{y}_{I},P_{a})\leq\mathcal{L}_{\mathbb{T}}(\mathbf{y}_{I},P_{b})\text{ and}\ a,b\in\{1,\cdots,n\}$
Here, $\mathbb{P}_{\mathbb{A}}(P|S)$ denotes the probability of the attacker LLM generating prompt $P$ conditioned on the reasoning string $S$. I.e., $\mathbb{P}_{\mathbb{A}}(P|S)=\prod_{i=1}^{l}\mathbb{P}_{\mathbb{A}}\big{(}p_{i}|[S,p_{1:i-1}]\big{)}$ where $P=[p_{1},\cdots,p_{l}]$. To evaluate this condition, we analyze the generation probabilities for a set of prompts before and after applying feedback to the reasoning string. When the attacker is Vicuna and the target is Llama-3-8B-RR, for 10 random tasks each with 10 iterations (totaling 100 function calls), we compute the difference in Cross-Entropy of the prompts conditioned on the original and updated reasoning strings: $-\log\Big{(}\mathbb{P}_{\mathbb{A}}\big{(}P_{a}|\mathbb{R}(S,F)\big{)}\Big{)}+\log\Big{(}\mathbb{P}_{\mathbb{A}}(P_{a}|S)\Big{)}$. If $a$ denotes the ordered index, this function must be increasing with respect to $a$ in order to satisfy Equation 5.1. Figure 4 presents the average results across all calls for $n=16$, showing that at each iteration, the curve roughly increases, with a slight decline in the last two prompts. This decline can be attributed to the "Lost in the Middle" effect, where models tend to focus more on the most recent text (Liu et al., 2023). Nonetheless, the value of the last two prompts do not fall below the value of the initial prompts, indicating that the generated feedback remains meaningful and not misleading. In Appendix C, We conduct two more experiments to shed more light on the functionality of Feedback LLM. As a sanity check, we demonstrate that the reversing the order of the given attacking prompts to the Feedback LLM leads to semantically reversed feedbacks, and a drop in the ASR due to moving in the wrong direction in the prompt space.
@@
The standard version of PAIR runs for only 3 iterations. We increased this number to $T=15$, matching it with our algorithm. Both algorithms use Mixtral to attack Llama-2-7B. Figures 5(a) and 5(b) depict the number of tasks that are jailbroken at each iteration for PAIR and our algorithm, respectively. PAIR achieved only two additional successful jailbreaks after the third iteration, accounting for 11% of successful attempts. In contrast, our algorithm accomplished 37% of jailbreaks after the third iteration, demonstrating a more effective utilization of iterations. Distributions of jailbreaks for other models are presented in Figures 8 and 9. As anticipated, a higher percentage of jailbreaks occur in later iterations for models with stricter safety measures that necessitate more extensive search; for instance, 57% after the third iteration for Llama-3-8B-RR (Figure 8(a)). PAIR relies on the reasoning capabilities of the attacker LLM to modify subsequent attacking prompts. Therefore, after encountering a few refusals from the target model, PAIR tends to deviate from the original intent to avoid the refusals. A couple of examples of this phenomenon are presented in Appendix E.